OpenStudy (anonymous):

The hypotenuse of a right triangle is 10 centimeters, and the shorter leg is 5 centimeters. Find the length of the other leg. A. 5√5 cm B. 5√3 cm C. 7 cm

OpenStudy (anonymous):

Pythagorean Theorem \[a^2+5^2=10^2\] \[a^2=100-25=75\] \[a=\sqrt{75}=5\sqrt{3}\]